OpenStudy (anonymous):

the perimeter of a rectangle is 138inches. the length exceeds the width by 7 inches. find the length and width

myininaya (myininaya):

P=2W+2L=138 L=7+W

OpenStudy (anonymous):

2l+2b=138 l-b=7 l=38 b=31